analyze the convergence/divergence of the next seriesυ) Σ* (ο <h <) 2=1 24 1) X vii) i) Σ η = 1 31 (α > 1, k 50 ) η=1 u?
Answer #1

8 W (17q xo) n=1 pn.qn ana pran (pan Ontl = (eaght (Panti appley ratio test dem / 9n+ / - him cpanel ah (pan hem h lea, pa Sewhen pq=1 then & ton E (1) n-1 pastial sum-sou n K=1 partial sum es unboundled then then series es divesgent. 1.0 series es c1 nk an (Kyo, as, n=1 an (ntik Conti anti ana (n+19k an + 1k Antl un ana АК a. nk a á + 1)* == (1+5)* Monte ) = allt)* à (tok